A fruit salad is made up of canteloupe, strawberries and banannas. If a salad weighing 135 oz. contains 3oz of banannas for every 1oz of strawberries and has the same number of oz. of cantaloupe as strawberries, how many oz. of banannas are in the salad?

Thanks!

Answer by ptaylor(2198)


Let x=amount of strawberries
Then x= amount of cantaloupes also
And 3x=amount of bananas
strawberries + cantaloupes + bananas = 135 oz. So our equation to solve:
x+x+3x=135
5x=135 divide both sides by 5
x=27 oz of strawberries and 27 oz of cantaloupe
3x=3*27=81 oz -----------------------------of bananas
ck
27+27+81=135
135=135

Hope this helps----ptaylor